OpenStudy (anonymous):

Use the given conditions to write an equation for the line in point slope form, slope intercept form and standard form. Passing through the points (5,-6) and perpendicular to the line whose equation is 3x-2y=9

OpenStudy (anonymous):

original slope y=3/2 x -9/2 slope = 3/2 perpendicular line's slope is -2/3 point slope y+6=(-2/3)(x-5) slope intercept y = (-2/3)x-(8/3) standard 2x+3y=-8